# Equality Saturation for Optimizing High-Level Julia IR

### v1

@@ 6. Limitations
As discussed in Section 3.2, our approach to converting IR to an e-graph using the CFG skeleton currently does not allow applying rewrites that alter control flow. Similarly, rewrite opportunities where part of an expression is hidden behind a function call are in general not considered. We optimize each function separately, applying rewrite rules both before and after inlining is applied. Other intermediate representations, discussed further in Section 7, could be used to represent the entire program in an e-graph and allow rewriting across call boundaries and control flow.
@@
In this work, we have proposed an ILP formulation for optimal e-graph extraction. The problem of extracting e-nodes from a general e-graph is NP-hard, which means that there is no guarantee that an optimal solution can be found (Goharshady et al., 2024). Currently, when the ILP solver is unable to find a solution, we fall back to a greedy extraction algorithm that is built into Metatheory.jl. In this work, we chose a simple cost model that treats the cost of nodes independent of other nodes. A more complex cost model could take into account more complex behavior that depends on multiple nodes at once.
@@
As discussed in Section 3, our rule syntax takes one Julia expression and transforms it into another. This is an intuitive syntax for expressing rewrite rules, but does not fully cover all the possible rules a developer could possibly want to express. For one, the rewrite rules can only be used to rewrite a single expression into another expression. Multi-pattern rewrite rules to match against and to rewrite multiple expressions at once are currently not supported. We also do not yet support rewrite rules matching against functions calls taking an unknown, variable amount of variables. This means that in certain cases, a developer has to rewrite multiple similar rules to match the same function with different numbers of arguments.
@@ 7.1. Equality Saturation
Equality saturation and its use as a code optimization tool was first introduced by Tate et al. (Tate et al., 2009), where it was used to optimize Java bytecode. Willsey et al. (Willsey et al., 2021) introduced algorithmic improvements for equality saturation providing asymptotic speed-ups. This has led to an explosion of new work exploring different applications for equality saturation, from optimizing tensor programs (Wang et al., 2020; Yang et al., 2021; Smith et al., 2021), to compiler verification (Stepp et al., 2011; Kourta et al., 2022), optimizations for specialized hardware (Ustun et al., 2022; VanHattum et al., 2021; Matsumura et al., 2023), and other purposes (Panchekha et al., 2015; Chandrakana Nandi et al., 2021).
Metatheory.jl (Cheli, 2021b) provides a framework for equality saturation in Julia. It has already been used for code optimization on symbolic representations of Julia code to speed up PDE solvers (Gowda et al., 2022). That work, however, only handles pure, symbolic programs, as it operates on a symbolic representation of Julia code obtained through manual expression building or tracing through heavily restricted, straight-line code. In contrast to our work, that existing use of Metatheory.jl does hence not allow the presence of general control flow or side effects.
To the best of our knowledge, by introducing type propagation in the e-graph, we are the first to apply equality saturation to a dynamic programming language. This allows rewrites to coexist and complement Julia’s multiple dispatch feature, as discussed in Section 4.
@@ 7.2. Intermediate Representations and E-Graphs
Cranelift, a compiler back-end for WebAssembly and Rust, is among the first to use e-graphs and equality saturation in a production-grade compiler (Fallin, 2023). While we reuse ideas and methods first introduced in the Cranelift compiler, we implemented our framework as a standalone project, not reusing any Cranelift components. Our code is integrated in the Julia compiler using user-land compiler extensions.
The Cranelift compiler does not do full equality saturation, but only recognizes equalities the first time a node is inserted in the e-graph. This implies that it is possible for equalities to remain undiscovered, but it obviates the fixed-point loop present in full equality saturation and ensures there are no cycles in the e-graph which allows for a more efficient representation of the graph in memory and leads to easier extraction. We use the same e-graph representation of code as used in Cranelift, but by utilizing an extraction scheme that enforces acyclic extraction, the optimizer is able to carry out full equality saturation. Whereas Cranelift and our work operate on IR that is complemented by a CFG, other work has sought to find alternative IR formats that allow representing control flow fully in the e-graph. Tate et al. (Tate et al., 2009) have introduced Program Expression Graphs (PEGs) which represent constructs such as loops as special expressions and show that PEGs can be used to do equality saturation. Similarly, Regionalized Value State Dependency Graphs (RVSDGs) (Bahmann et al., 2015; Reissmann et al., 2020) represent control flow as (nested) expressions. A research prototype already exists using RVSDGs in conjunction with equality saturation for a simple toy language (Sharp, 2022). Our approach remains closer to the original SSA-based IR of the source language, allowing less expensive conversion routines.
Willsey et al. introduced e-class analyses to associate additional metadata with each e-class (Willsey et al., 2021). The metadata is kept up to date throughout the equality saturation process by potential modifications whenever new information is available, for example, when e-classes are merged. An exemplary existing use case for an e-class analysis is constant propagation, where each e-class can optionally carry a constant value. To track types in the e-graph, we designed a novel e-class analysis, as discussed in Section 3.3
@@ 7.3. Extraction
We based our ILP optimal extraction formulation on the work of He et al. (He et al., 2017). Recent work proposes other formulations, for example by looking at e-graphs through the lens of circuits (Sun et al., 2024), or finite state automata (Y. Wang et al., 2022). These formulations can lead to faster extraction or provide termination guarantees under particular assumptions. Cranelift operates in a Just-In-Time context where fast and reliable extraction is crucial. For this reason, they use an extraction algorithm that greedily tries to minimize the total number of nodes, without taking into account node reuse in extracted expressions or from expressions that dominate them. In contrast to Cranelift’s greedy algorithm, we have introduced an ILP formulation that aims for node reuse from other nodes within an extracted expression, as well as from nodes in expressions that dominate them in the CFG.
@@ 7.4. Julia Code Optimization
Other works exist that aim to improve the performance of Julia code with custom optimizations. Symbolics.jl (Gowda et al., 2022) can be used to transform scientific machine learning code into a symbolic representation at which point simplifications can be applied. Finch.jl (Ahrens et al., 2024, 2023) uses a custom IR to represent iteration patterns of loop nests over sparse or structured arrays. We instead focus on optimizing code in Julia’s own IR format, allowing users of our system to target any Julia code.
@@ 8. Conclusion and Future Work
We have developed a novel system that allows Julia developers to write domain-specific rewrite rules that are automatically applied using equality saturation and that works in the presence of control flow and side effects. Through an e-class analysis, we keep track of the most specific type of all equivalent terms and use this information to support type-constrained rewrite rules. Unlike previous work, our system can rewrite dynamically-typed code. We introduce a technique called CFG skeleton relaxation that allows rewrite rules to nullify side effects in the original code. We have adapted an ILP formulation for optimal, acyclic e-graph extraction to take into account value reuse from dominating statements. Finally, we show how this system can perform rewrites on high-level domain-specific code and how the use of equality saturation solves the phase ordering problem.
In the future, our work can be extended to support rewriting control flow and apply rewrites that span different call depths. Further improvements to our extraction scheme could also lead to more efficient code being generated.

### v2

@@ 7. Limitations
As discussed in Section 3, our approach to converting IR to an e-graph using the CFG skeleton currently does not support rewrites that alter control flow. Similarly, since we apply rewrites on code where function calls have not yet been inlined, rewrite opportunities where part of an expression is hidden behind a function call are not considered. We optimize each method separately. Other intermediate representations, discussed in Section 8, could be used to represent the entire program in an e-graph and allow rewriting across call boundaries and control flow.
@@
In this work, we have proposed an ILP formulation for optimal e-graph extraction. The problem of extracting e-nodes from a general e-graph is NP-hard (Goharshady et al., 2024), which means that there is no guarantee that an optimal solution can be found in polynomial time. Currently, when the ILP solver is unable to find a solution within a given timeframe, we fall back to a greedy extraction algorithm that is built into Metatheory.jl. In this work, we chose a simple cost model that treats the cost of nodes independent of other nodes. A more complex cost model could take into account more complex behavior that depends on multiple nodes at once.
@@
As discussed in Section 3, our rule syntax takes one Julia expression and transforms it into another. This is an intuitive syntax, but does not fully cover all the possible rules a developer could possibly want to express. For one, the rewrite rules can only be used to rewrite a single expression into another expression. Multi-pattern rewrite rules to match against and to rewrite multiple expressions at once are currently not supported. We also do not yet support rewrite rules matching against function calls taking an unknown, variable amount of variables. A developer may hence have to rewrite multiple similar rules for the same function with different argument counts.
Rule applicability
Section 6 shows that certain rewrites do not always yield positive returns. Consequently, the burden is currently on the programmer to select the rule sets for their particular problem. In the future, this could be automated using an improved cost model, in combination with more detailed program analysis to gauge the effectiveness of rewrite rules. Alternatively, the rule applicability could be made more extensible, allowing package developers to decide programmatically when a rule can be applied.
@@ 8.1. Equality Saturation
EqSat and its use as a code optimization tool was first introduced by Tate et al. (Tate et al., 2009), where it was used to optimize Java bytecode. Willsey et al. (Willsey et al., 2021) introduced algorithmic improvements for EqSat providing asymptotic speed-ups. This has led to an explosion of new work exploring different applications for EqSat, from optimizing tensor programs (Wang et al., 2020; Yang et al., 2021; Smith et al., 2021), to compiler verification (Stepp et al., 2011; Kourta et al., 2022), optimizations for specialized hardware (Ustun et al., 2022; VanHattum et al., 2021; Matsumura et al., 2023), and other purposes (Panchekha et al., 2015; Nandi et al., 2021).
Metatheory.jl (Cheli, 2021b) provides a framework for EqSat in Julia. It has already been used for code optimization on symbolic representations of Julia code to speed up PDE solvers (Gowda et al., 2022). That work, however, only handles pure, symbolic programs, as it operates on a symbolic representation of Julia code obtained through manual expression building or tracing through heavily restricted, straight-line code. In contrast to our work, the existing use of Metatheory.jl does hence not allow the presence of general control flow or side effects.
To the best of our knowledge, by introducing type propagation in the e-graph, we are the first to apply EqSat to a dynamic programming language. This allows rewrites to coexist and complement Julia’s multiple dispatch feature, as discussed in Section 4.
@@ 8.2. Intermediate Representations and E-Graphs
Cranelift, a compiler back-end for WebAssembly and Rust, is among the first to use e-graphs and EqSat in a production-grade compiler (Fallin, 2023). While we reuse ideas and methods first introduced in the Cranelift compiler, we implemented our framework as a standalone project, not reusing any Cranelift components. Our code is integrated in the Julia compiler using user-land compiler extensions.
The Cranelift compiler does not do full EqSat, but only recognizes equalities the first time a node is inserted in the e-graph. This implies that it is possible for equalities to remain undiscovered, but it obviates the fixed-point loop present in full EqSat and ensures there are no cycles in the e-graph which allows for a more efficient representation of the graph in memory and leads to easier extraction. We use the same e-graph representation of code as used in Cranelift, but by utilizing an extraction scheme that enforces acyclic extraction, the optimizer is able to carry out full EqSat. Whereas Cranelift and our work operate on IR that is complemented by a CFG, other work has sought to find alternative IR formats that allow representing control flow fully in the e-graph. Tate et al. (Tate et al., 2009) have introduced Program Expression Graphs (PEGs) which represent constructs such as loops as special expressions and show that PEGs can be used to do EqSat. Laird et al. (Laird et al., 2024) raise supported LLVM IR into a functional IR that also allows control flow transformations. Similarly, Regionalized Value State Dependency Graphs (RVSDGs) (Bahmann et al., 2015; Reissmann et al., 2020) represent control flow as (nested) expressions. A research prototype already exists using RVSDGs in conjunction with EqSat for a simple toy language (Sharp, 2022). Our approach remains closer to the original SSA-based IR of the source language, allowing less expensive conversion routines. Lastly, there are multiple works applying equality saturation on MLIR code (Merckx et al., 2025; Zayed and Dubach, 2025). These do not keep track of side effects and control flow in a CFG skeleton and thus can only operate on code that is pure. Additionally, they operate on statically typed IR in contrast to our work.
Willsey et al. introduced e-class analyses, a framework to associate additional metadata with each e-class (Willsey et al., 2021). The metadata is kept up to date throughout the EqSat process by potential modifications whenever new information is available, for example, when e-classes are merged. An exemplary existing use case for an e-class analysis is constant propagation, where each e-class can optionally carry a constant value. Further work has shown that this technique can be combined with abstract interpretation for more complex program analyses (Coward et al., 2023b). To track types in the e-graph, we designed a novel e-class analysis, as discussed in Section 3.2.
@@ 8.3. Extraction
We based our ILP optimal extraction formulation on the work of He et al. (He et al., 2023). Recent work proposes other formulations, for example by looking at e-graphs through the lens of circuits (Sun et al., 2024), or finite state automata (Y. Wang et al., 2022). These formulations can lead to faster extraction or provide termination guarantees under particular assumptions. Cranelift operates in a just-in-time context where fast and reliable extraction is crucial. For this reason, they use an extraction algorithm that greedily tries to minimize the total number of nodes, without taking into account node reuse in extracted expressions or from expressions that dominate them. In contrast to Cranelift’s greedy algorithm, we have introduced an ILP formulation that aims for node reuse from other nodes within an extracted expression, as well as from nodes in expressions that dominate them in the CFG.
@@ 8.4. Julia Code Optimization
Other works exist that aim to improve the performance of Julia code with custom optimizations. Symbolics.jl (Gowda et al., 2022) can be used to transform scientific machine learning code into a symbolic representation on which simplifications can be applied. Finch.jl (Ahrens et al., 2024, 2023) uses a custom IR to represent iteration patterns of loop nests over sparse or structured arrays. We instead focus on optimizing code in Julia’s own IR format, allowing users of our system to target any Julia code.
@@ 9. Conclusion and Future Work
We have developed a novel system that allows Julia developers to write domain-specific rewrite rules that are automatically applied using equality saturation. Our system works in the presence of control flow and side effects. Through an e-class analysis, we keep track of the most specific type of all equivalent terms and use this information to support type-constrained rewrite rules. Unlike previous work, our system can rewrite dynamically-typed code. We introduce a technique called CFG skeleton relaxation that allows rewrite rules to nullify side effects in the original code. We have adapted an ILP formulation for optimal, acyclic e-graph extraction to take into account value reuse from dominating statements. Finally, we show how this system can perform rewrites on high-level domain-specific code and how the use of equality saturation solves the phase-ordering problem.
In the future, our work can be extended to support rewriting control flow and apply rewrites that span different call depths, by integrating function call inlining in the equality saturation procedure. Further improvements to our extraction scheme could also lead to more efficient code being generated.